OpenStudy (anonymous):

Given the system of equations, what is the value of the system determinant? x + y = 8 x - y = 10 0 -1 -2

OpenStudy (unklerhaukus):

\[\left|\begin{array} \\1 & 1 \\ 1 & -1 \\ \end{array}\right|=(1\times-1)-1\times1=-1-1=-2\]

OpenStudy (anonymous):

-1

OpenStudy (unklerhaukus):

/?

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

how did you get -1

OpenStudy (anonymous):

ya

OpenStudy (unklerhaukus):

how

OpenStudy (anonymous):

times

OpenStudy (anonymous):

\[\left[\begin{matrix}a & b \\ c& d\end{matrix}\right]\] = a.d - b.c

OpenStudy (anonymous):

ITS 1